How do you write the standard form of an equation of the line that passes through each pair of points? (6,5) (12,-3)
Answer by mananth(16946)

x1 y1 x2 y2
6 5 12 -3

slope m = (y2-y1)/(x2-x1)
( -3 - 5 )/( 12 - 6 )
( -8 / 6 )
m= -4/3
The required line passes through both the points. So we select one point and plug it with the slope
Plug value of the slope and point ( 6 , 5 ) in
Y = m x + b
5.00 = -4/3 * 6 + b
5.00 = -8 + b
b= 5 + 8
b= 13
So the equation will be
Y = -4/3 x + 13
multiply equation by 3
3y=-4x+39
4x+3y=39 is the equation

m.ananth@hotmail.ca